# revtex4-2Repair the float

### v1

@@ C.3.1 Relation to usual definitions in isotropic materials
where in the second to last equality we used $i^{2}=-1$ and $\varepsilon_{lim}=-\varepsilon_{ilm}$. Now, for $k=l$ we have $\varepsilon_{jli}\varepsilon_{ilm}j_{jm}^{(0)}=-\varepsilon_{jli}\varepsilon_{mli}j_{jm}^{(0)}=\sum_{m\neq l}j_{mm}^{(0)}$, while for $k\neq l$ we have $\varepsilon_{jki}\varepsilon_{ilm}j_{jm}^{(0)}=\varepsilon_{jki}\varepsilon_{lmi}j_{jm}^{(0)}=j_{mk}^{(0)}$. Hence we may write the contribution of the spin-swapping term as $j_{mk}^{(0)}-\delta_{kl}j_{mm}^{(0)}$. With this we may express the current as
@@
that is, $\theta$ and $\varkappa$ are the spin-Hall and spin-swapping coefficients consistent with [58, 59].
@@ Appendix D Vortex in a gyrotropic superconductor
In this section we study the excess spin around a vortex in a gyrotropic superconductor. The pair amplitudes vanish at the center ($r=0$), while the phase winds as $e^{in\phi}$ around the center of the vortex, where $\phi$ denotes the tangential coordinate and $n$ is the winding number of the vortex. Such vortices can be created for example in a type II superconductor by applying a magnetic field. If the magnetic field is subsequently turned off, the vortices persist for a finite time, since they can only disappear via annihilation with a vortex with opposite winding or via the boundaries.
Now, according to Eq. (11) in the main body, a gradient in the Green’s function is coupled to the spin via the torque $\gamma_{kj}\sigma_{j}\partial_{k}g_{s}$, where $g_{s}$ is the Green’s function in the absence of a spin-galvanic effect. We may distinguish two different kinds of spatial dependence, spatial dependence of the phase and spatial dependence of the magnitude. Indeed, for a general system we may write
$g_{s}(\bm{r})=e^{-i\phi(\bm{r})\tau_{3}/2}(g_{s0}(\bm{r})\tau_{3}+f_{s0}\bm{r}\tau_{2})e^{i\phi(\bm{r})\tau_{3}/2},$ (60)
If the phase depends on position, the spatial derivative has opposite sign for the $\langle\psi\psi\rangle$ and $\langle\psi^{\dagger}\psi^{\dagger}\rangle$ correlations, and it appears as $\nabla\phi[\tau_{3},g]$. Thus, in the presence of a spin-galvanic term the Usadel equation obtains an extra term $\partial_{k}\phi\gamma_{lk}[\sigma_{l}\tau_{3},g]$. This term appears exactly like an exchange field.
On the other hand, if the magnitude of the pair amplitudes depends on position, there is no time-reversal symmetry breaking, and hence it is not similar to an exchange field, it can not introduce any spin dependence in the density of states. Indeed, considering that the normalization condition needs to be satisfied, but the phase is not affected, such terms are proportional to $[e^{i\phi\tau_{3}}\tau_{2},g]$, and we conclude it is merely a source of triplet correlations.
@@
Writing the Usadel equation of the system with the vortex in Riccati parameterization, using the notation $a,b$ instead of the usual $\gamma,\tilde{\gamma}$ to avoid confusion with the tensor of the gyrotropic term,
$g=\begin{bmatrix}(1+ab)^{-1}(1-ab)\text{sign}(\omega_{n})&2(1+ab)^{-1}a\\ 2(1+ba)^{-1}b&-(1+ba)^{-1}(1-ba)\text{sign}(\omega_{n}),\end{bmatrix}$ (62)
the equations can be found by substituting the Riccati parameterization in a term $A$ and then calculating the contribution to the equation for $\nabla^{2}a$ via $\frac{1}{4}(\mathbf{1}+ab)\Big{(}\text{sign}(\omega)\text{Tr}\sigma_{d}(\tau_{1}+i\tau_{2})A-\text{Tr}\sigma_{d}(\mathbf{1}+\tau_{3})Aa\Big{)}$, where $d=0,1,2,3$, while the contribution to the equation for $b$ can be found via $\frac{1}{4}(\mathbf{1}+ba)\Big{(}-\text{sign}(\omega)\text{Tr}\sigma_{d}(\tau_{1}-i\tau_{2})A-\text{Tr}\sigma_{d}(\mathbf{1}-\tau_{3})Ab\Big{)}$.
For a single vortex, in the absence of a gyrotropic tensor the equation for $a$ can be written as
@@
while the equation for $b$ is found similarly. In polar coordinates this becomes
@@
$I(r)=\frac{-i}{1+ab}(a\nabla b-b\nabla a)=\frac{1}{r}\frac{na_{s}^{2}}{1+a_{s}^{2}}\bm{\hat{\phi}}.$ (69)
@@
$=\frac{2|\omega|}{D}\delta a_{t}+\frac{2\Delta}{D}a_{s}\delta b_{t}+\frac{\Delta}{D}(\delta b_{t}e^{in\phi}-\delta a_{t}e^{-in\phi})+\gamma_{o}(\frac{in}{r}a_{s}\sigma_{x}e^{i\phi\sigma_{z}}-\frac{\partial a_{s}}{\partial r}\sigma_{y}e^{i\phi\sigma_{z}})e^{in\phi}$ (70)
@@
$=\frac{2|\omega|}{D}\delta b_{t}+\frac{\Delta}{D}(2a_{s}\delta a_{t}-(\delta b_{t}e^{in\phi}-\delta a_{t}e^{-in\phi}))-\gamma_{o}(-\frac{in}{r}a_{s}\sigma_{x}e^{i\phi\sigma_{z}}+\frac{\partial a_{s}}{\partial r}\sigma_{y}e^{i\phi\sigma_{z}})e^{-in\phi}$ (71)
@@
$=\frac{2|\omega|}{D}a_{x}+\frac{1}{l_{\text{so}}^{2}}a_{x}+\frac{2\Delta}{D}a_{s}a_{x}+\frac{2\Delta}{D}\text{Re}a_{x}+\frac{in}{r}\gamma_{o}a_{s}$ (72)
@@
$=\frac{2|\omega|}{D}a_{y}+\frac{1}{l_{\text{so}}^{2}}a_{y}+\frac{2\Delta}{D}a_{s}a_{y}+\frac{2\Delta}{D}\text{Re}a_{y}-\gamma_{o}\frac{\partial a_{s}}{\partial_{r}}$ (73)
@@
$=\frac{2|\omega|}{D}a_{1}+\frac{1}{l_{\text{so}}^{2}}a_{1}+\frac{2\Delta}{D}a_{s}a_{1}+\frac{n}{r}\gamma_{o}a_{s}$ (74)
@@
$=\frac{2|\omega|}{D}a_{2}+\frac{1}{l_{\text{so}}^{2}}a_{2}+\frac{2\Delta}{D}(1+a_{s})a_{2}-\gamma_{o}\frac{\partial a_{s}}{\partial_{r}}$ (75)
@@
The results are shown in Fig. 1. We show the current density $I$ normalized by the maximum current density and the excess spin density $S^{l}$ normalized by the maximum, as a function of the radial coordinate, which is normalized by $\xi=\sqrt{D/2\Delta(T)}$. Clearly, for low temperatures, the spatial dependence of the excess spin density differs from the spatial dependence of the current density, with the spin density profile having its maximum further away from the vortex core. This is in contrast with the GL theory prediction, in which the spin depends linearly on the local current. For larger temperatures the difference becomes smaller. However, we notice that even for $T/T_{c}=0.98$, i.e. close to the critical current, there is still a noticeable difference between the two spatial profiles.
Figure 3: The current density ($I$) and excess spin density ($S$) around a vortex with $n=1$ normalized to their maximum values as a function of the radial coordinate. The maximum in the excess spin density is further away from the vortex core than the maximum of the current density. The difference becomes smaller as the temperature approaches the critical temperature.
Appendix E Ginzburg-Landau
In this section we derive the Ginzburg-Landau (GL) free energy from the Usadel equation. The GL functional is a convenient description for superconductors at temperatures close to the critical temperature, so that $\Delta(T)\ll k_{B}T$. Therefore, near the critical temperature, we may write the Usadel equation expanded in pair amplitudes and use the solution in the pair amplitudes combined with the self-consistency relation to obtain the GL functional as has been shown before for both the linearized Usadel and Eilenberger equations [60, 61, 62, 9]. We will here follow an approach similar to the one presented in those articles. We approximate the quasiclassical Green’s function as
@@ Appendix E Ginzburg-Landau
$F_{Z}=\frac{1}{2}\chi_{ij}\tilde{h}_{i}\tilde{h}_{j}$ (78)
where $\chi_{ij}$ is the response tensor. The Lifshitz-invariant is the term proportional to $q$ of the previous expression. Thus,

### v2

@@ C.3.1 Isotropic materials
where in the second to last equality we used $i^{2}=-1$ and $\varepsilon_{lim}=-\varepsilon_{ilm}$. For $k=l$ we have $\varepsilon_{jli}\varepsilon_{ilm}j_{jm}^{(0)}=-\varepsilon_{jli}\varepsilon_{mli}j_{jm}^{(0)}=\sum_{m\neq l}j_{mm}^{(0)}$. On the other hand, for $k\neq l$ we have $\varepsilon_{jki}\varepsilon_{ilm}j_{jm}^{(0)}=\varepsilon_{jki}\varepsilon_{lmi}j_{jm}^{(0)}=j_{mk}^{(0)}$. Hence we may write the contribution of the spin-swapping term as $j_{mk}^{(0)}-\delta_{kl}j_{mm}^{(0)}$. With this we may express the current as
@@
that is, $\theta$ and $\varkappa$ are the spin-Hall and spin-swapping coefficients in isotropic systems consistent with previous works [58, 59].
@@ Appendix D Vortex in a gyrotropic superconductor
In this section we study the excess spin around a vortex in a gyrotropic superconductor. For a vortex, the pair amplitudes vanish at the center ($r=0$), while the phase of the pair amplitudes winds as $e^{in\phi}$ around the center of the vortex, where $\phi$ denotes the tangential coordinate and $n$ is the winding number of the vortex. Such vortices can be created for example in a type II superconductor by applying a magnetic field.
Now, according to Eq. (11) in the main body, a gradient in the Green’s function is coupled to the spin via the term in the torque $\gamma_{kj}\sigma_{j}\partial_{k}g_{s}$, where $g_{s}$ is the Green’s function in the absence of a spin-galvanic effect. We may distinguish two different kinds of spatial dependence, spatial dependence of the phase and spatial dependence of the magnitude. Indeed, for a general system, we may write
$g_{s}(\bm{r})=e^{-i\phi(\bm{r})\tau_{3}/2}(g_{s0}(\bm{r})\tau_{3}+f_{s0}(\bm{r})\tau_{2})e^{i\phi(\bm{r})\tau_{3}/2},$ (60)
If the phase depends on position, the spatial derivative has opposite sign for the $\langle\psi\psi\rangle$ and $\langle\psi^{\dagger}\psi^{\dagger}\rangle$ correlations, and it appears as $\nabla\phi[\tau_{3},g]$. Thus, in the presence of a spin-galvanic term the Usadel equation has an additional term $\partial_{k}\phi\gamma_{lk}[\sigma_{l}\tau_{3},g]$. This term appears exactly like an exchange field.
On the other hand, if the magnitude of the pair amplitudes depends on position, there is no time-reversal symmetry breaking, and hence it is not like an exchange field, and it can not introduce any spin dependence in the density of states.
@@
We now compute explicitly the current and excess spin densities by solving the Usadel equation, Eq. (7), assuming an isotropic spin-relaxation $\Gamma_{jk}=\frac{D}{l_{\text{so}^{2}}}\delta_{jk}$, where $l_{\text{so}}$ is the spin-relaxation length, for simplicity of notation. For this we use the Riccati parameterization (we use the notation $a,b$ instead of the usual $\gamma,\tilde{\gamma}$ to avoid confusion with the tensor of the gyrotropic term)
$g=\begin{bmatrix}(1+ab)^{-1}(1-ab)\text{sign}(\omega_{n})&2(1+ab)^{-1}a\\ 2(1+ba)^{-1}b&-(1+ba)^{-1}(1-ba)\text{sign}(\omega_{n})\end{bmatrix}\;.$ (62)
This parametrization guarantees the normalization condition $g^{2}=1$. For a single vortex, in the absence of a gyrotropic tensor the equation for $a$ can be written as (cf. Eq. (7))
@@
while the equation for $b$ is found similarly. In polar coordinates this equation can be written as
@@
$I(r)=\frac{-i}{1+a_{0}b_{0}}(a_{0}\nabla b_{0}-b_{0}\nabla a_{0})=\frac{1}{r}\frac{na_{s}^{2}}{1+a_{s}^{2}}\bm{\hat{\phi}}.$ (68)
@@
$=\frac{2|\omega|}{D}\delta a_{t}+\frac{2\Delta}{D}a_{s}\delta b_{t}+\frac{\Delta}{D}(\delta b_{t}e^{in\phi}-\delta a_{t}e^{-in\phi})+\gamma_{o}(\frac{in}{r}a_{s}\sigma_{x}e^{i\phi\sigma_{z}}-\frac{\partial a_{s}}{\partial r}\sigma_{y}e^{i\phi\sigma_{z}})e^{in\phi}\;,$ (69)
@@
$=\frac{2|\omega|}{D}\delta b_{t}+\frac{\Delta}{D}(2a_{s}\delta a_{t}-(\delta b_{t}e^{in\phi}-\delta a_{t}e^{-in\phi}))-\gamma_{o}(-\frac{in}{r}a_{s}\sigma_{x}e^{i\phi\sigma_{z}}+\frac{\partial a_{s}}{\partial r}\sigma_{y}e^{i\phi\sigma_{z}})e^{-in\phi}\;.$ (70)
@@
$=\frac{2|\omega|}{D}a_{x}+\frac{1}{l_{\text{so}}^{2}}a_{x}+\frac{2\Delta}{D}a_{s}a_{x}+\frac{2\Delta}{D}\text{Re}a_{x}+\frac{in}{r}\gamma_{o}a_{s}\;,$ (71)
@@
$=\frac{2|\omega|}{D}a_{y}+\frac{1}{l_{\text{so}}^{2}}a_{y}+\frac{2\Delta}{D}a_{s}a_{y}+\frac{2\Delta}{D}\text{Re}a_{y}-\gamma_{o}\frac{\partial a_{s}}{\partial_{r}}\;.$ (72)
@@
$=\frac{2|\omega|}{D}a_{1}+\frac{1}{l_{\text{so}}^{2}}a_{1}+\frac{2\Delta}{D}a_{s}a_{1}+\frac{n}{r}\gamma_{o}a_{s}\;,$ (73)
@@
$=\frac{2|\omega|}{D}a_{2}+\frac{1}{l_{\text{so}}^{2}}a_{2}+\frac{2\Delta}{D}(1+a_{s})a_{2}-\gamma_{o}\frac{\partial a_{s}}{\partial_{r}}\;.$ (74)
@@
The results are shown in Fig. 3, of which the left panel corresponds to Fig. 1 in the main text. We show the current density $I$ normalized by the maximum current density and the excess spin density $S^{l}$ normalized by the maximum, as a function of the radial coordinate, which is normalized by $\xi=\sqrt{D/2\Delta(T)}$. Clearly, for low temperatures, the spatial dependence of the excess spin density differs from the spatial dependence of the current density, with the spin density profile having its maximum further away from the vortex core. This is in contrast with the GL theory prediction, in which the spin depends linearly on the local current. For larger temperatures the difference becomes smaller. However, we notice that even for $T/T_{c}=0.98$, i.e. close to the critical current, there is still a noticeable difference between the two spatial profiles.
Figure 3: The current density ($I$) and excess spin density ($S$) around a vortex with $n=1$ normalized to their maximum values as a function of the radial coordinate, for (left) $T/T_{c}=0.05$ and (right) $T/T_{c}=0.98$. The maximum in the excess spin density is further away from the vortex core than the maximum of the current density. The difference becomes smaller as the temperature approaches the critical temperature.
Appendix E The Ginzburg-Landau limit: Derivation of the Lifshtiz-invariant
In this section we derive the Ginzburg-Landau (GL) free energy from the Usadel equation. The GL functional is a convenient description for superconductors at temperatures close to the critical temperature, so that $\Delta(T)\ll k_{B}T$. Therefore, near the critical temperature, we may linearize the Usadel equation in pair amplitudes and use the solution combined with the self-consistency relation to obtain the GL functional as has been shown before for both the linearized Usadel and Eilenberger equations [60, 61, 62, 9]. We will here follow an approach similar to the one presented in those articles. We approximate the quasiclassical Green’s function as
@@ Appendix E The Ginzburg-Landau limit: Derivation of the Lifshtiz-invariant
$F_{Z}=\frac{1}{2}\chi_{ij}\tilde{h}_{i}\tilde{h}_{j}\;,$ (77)
where $\chi_{ij}$ is the response tensor for the excess spin. The Lifshitz-invariant is the term proportional to $q$ of the previous expression. Thus,
@@
This expression is identical to the expression for the Lifshitz invariant of the Ginzburg-Landau functional presented in [49] we define the second rank pseudotensor as $K_{jk}=\chi_{ij}\gamma_{ik}/\Delta^{2}$. However, since a phase gradient appears in the same way as a magnetic field even in the nonlinear equation, is the lowest order response to $h_{k}+q_{l}\gamma_{kl}$ for any temperature has the form of Eq. (77), with suitable definition of $\chi_{ij}(T)$. Therefore, Eq. (78) is more generic than the expression for the Lifshitz invariant in [49], it represents the contribution to the Free energy linear in the phase gradient for any temperature.